# Designer three-dimensional electronic bands in asymmetric transition metal dichalcogenide heterostructures

### v1

@@
To demonstrate how this concept of achieving bulk-like band characteristics with mixed material properties can be realised in real systems, Figure 2 overviews the electronic structure of a (WSe2)3/(MoSe2)5 (Fig. 2(a)) and a (MoSe2)3/(WSe2)3 (Fig. 2(e)) heterostructure constructed by interfacing two few-layer TMD flakes on graphite/n-type silicon substrates (see Methods). This pair of heterostructures thus closely approximate opposite surfaces of a single MoSe2-WSe2 system. Photoemission spectra acquired from these heterostructure regions are displayed in Fig. 2(b, f).
@@
As exemplified in Fig. 1(b-e), there is a one-to-one correspondence between the number of discrete $k_{z}$ sub-bands within the $d_{z^{2}}$-manifold at the $\Gamma$ point of 2H-structured TMDs and the number of monolayers within the system, Chang et al. (2014); Le Févre et al. (2024). However, as basic band theory stipulates that periodicity in momentum space cannot be achieved without periodicity in real space, the number of discrete $k_{z}$ sub-bands visible within the stacked regions would not be expected to reflect the full extent of the heterostructures which lack a well-defined unit cell. Yet, the photoemission spectra in Fig. 2(b,f) demonstrate that the number of sub-bands is not only increased relative to equivalent datasets acquired from the constituent flakes (shown in full in Supplemental Fig. 1), but also matches the total layer number in real space (i.e. eight for (b) and six for (f)) of the entire heterostructure. The electron mean free path of less than 0.7 nm, applicable to our photoemission experiments, precludes that the experimental result originates from a superposition of spectra from weakly-interacting top and bottom TMD flakes which each have thicknesses of $\approx$ 0.7 nm per monolayer Pan and Zahn (2022). This alternate scenario is further discarded through constant-energy $k_{x}$-$k_{y}$ contour mapping in Supplemental Figure 2 sup . Instead, the six or eight distinct $d_{z^{2}}$ sub-bands at $\Gamma$ are delocalised throughout the structure, and are therefore accessible to a surface sensitive probe.
@@
While counterintuitive, these findings are entirely consistent with the discussion surrounding Figure 1: The array of hybridized bands at $\Gamma$ belonging to the collective band structure of the finite six or eight layer crystal structure closely mimics the discretised $k_{z}$-projected band manifold of ordered systems. For bands derived from orbitals with significant interlayer hopping potential, the $k_{z}$ dispersion is therefore effectively constructed without the need for a periodic repeating unit. To further test this description, the experimental band dispersions of both WSe2-MoSe2 heterostructures are compared to layer-projected, density functional theory (DFT)-based slab calculations in Fig. 2(c,g), (see Methods) with zero twist angle. The fine structure around the $\Gamma$ point is extremely well reproduced in each case, verifying the present interpretation while also demonstrating both that collective electron phases not captured by these calculations, and the large twist angles chosen experimentally in order to decouple Moiré effects (Sup. Fig. 3), are not conditional or consequential for the formation of these delocalised states.
@@
To further scrutinize the observed three-dimensional behavior, in Figure 2(d) and (h) we track the evolution of the delocalised states at $\Gamma$ with changing photon energy, thus mapping their behavior with changing $k_{z}$ Damascelli (2004). These spectra can be compared to equivalent datasets for a three-layer MoSe2 and WSe2 flake (Sup. Fig. 1 (f,i)). For both the individual TMD flakes and the artificial heterostructures, each $d_{z^{2}}$-subband is two-dimensional (i.e. non-dispersive along $k_{z}$) in-line with their assignment as dimensionally confined states, though there is a photon energy dependence to their spectral weight. In the six and eight-layer hybrid systems (Fig. 2(d,h)), it becomes apparent that this shifting spectral weight between sub-bands is approximating the periodic $k_{z}$ dispersion that would be expected for a bulk 2H-structured TMD system. Indeed, the overlaid solid lines are a tight-binding model for an idealized bulk structure with alternating WSe2 and MoSe2 layers along the $c$-axis, once again showing how the periodic $k_{z}$-dispersions expected for ordered solids are emulated by interfacing pairs of TMDs in the intermediate layer regime.
[…]
The observed $c$-axis delocalisation, persisting even for this energetically-mismatched scenario, is facilitated by the lowering of the total energy of the system by partially lifting the constraints of 2D quantization. Along the $c$-axis, electrons form standing waves due to the finite thickness of the artificial crystals. This effect can be expected to vanish as one or both flakes approach the infinite-layer limit and electrons at the interface gain access to a bulk reservoir, ultimately preserving the quantized electronic structure of the remaining few-layer component. This is tested and shown explicitly in Supplemental Figure 5 through the characterization of a second overlap region within the sample in Fig. 2(i); between a 4ML-WSe2 flake atop a many-layer NbSe2 crystal. While charge transfer to the semiconducting top-layer still occurs in addition to some normalization of band features relative to the isolated 4ML-WSe2 top-layer, there are no additional well-defined $d_{z^{2}}$-subbands.
@@
While the presence of delocalised electron behavior within the $d_{z^{2}}$ orbital manifolds of the finite heterostructures is clear, the same is not true for regions of the electronic structure derived from in-plane orbitals. The inherent asymmetry to heterostructures constructed by interfacing two dissimilar TMD flakes ensures a spatial dependence to the broader electronic structure along the $c$-axis between limits defined by the distinct component materials. The scale of this dependence is dependent on orbital type. In Figure 3 and Supplemental Figs. 6-7, this real space evolution is explored in full. In Fig. 3(a), the orbitally-resolved WSe2 and MoSe2 terminations of the experimentally realized (MoSe2)3/(WSe2)3 heterostructure are compared. The valence band maxima at K and $\Gamma$, derived exclusively from $d_{\parallel}$ and $d_{z^{2}}$-orbitals respectively, represent the two extremes of $c$-axis delocalisation: The details of the spin-orbit splitting at K is determined only by the local crystal structure, while the band structure at $\Gamma$ is equivalent between terminations and indeed between the top and bottom layers (shown explicitly in Supplemental Figure 6). The rest of the band structure behaves similarly, with the real-space electronic structure dependence directly determined by the $d_{\parallel}$/ $d_{z^{2}}$ orbital ratio, further influenced by hybridization between bands of differing orbital origins.
@@
The asymmetry of our artificial heterostructures can be expected to modify the spin structure relative to natural TMD systems. As discussed earlier, in a natural 2H-TMD, each constituent TMD layer, owing to its $D_{3h}$ symmetry, can only exhibit an out-of-plane spin-orbit field, whose direction is opposite to that of the neighbouring layers. Consequently, this implies that in-plane spin textures, including chiral spin-momentum locking types arising from the Rashba effect Bychkov and Rashba (1984), are effectively suppressed. Furthermore, the hidden out-of-plane spin textures of each layer can only become accessible at the surface, where translational symmetry breaking disrupts the natural cancellation of $B_{\text{SO}}$ along the stacking axis.
@@
In contrast, our artificial structures can naturally accommodate both in-plane Rashba-type and out-of-plane Zeeman-type spin textures due to their unique heterostructural properties. Specifically, the chiral Rashba spin polarisation can emerge naturally from the potential gradient formed at the MoSe2/WSe2 interface and propagate throughout the entire structure via $d_{z^{2}}$ and other out-of-plane orbitals. To elucidate the implications of this, in Fig. 3(b-f), we compare a (MoSe2)3/(WSe2)3 heterostructure with an idealised six-layer (MoSe2/WSe2)3 system. The calculated layer-projected spin polarisation for the chiral ($x$) and out-of-plane ($z$) components is shown for both the asymmetric and idealised cases. As can be observed, in both configurations, strikingly similar patterns of in-plane and out-of-plane spin polarisation emerge, which remain valley-specific. Particularly, in both cases, the chiral ($x$) component is significantly enhanced around the $\Gamma$ point, where the $d_{z^{2}}$ contribution is dominant. As with the discussions surrounding the $k_{z}$ dispersions on show in Fig. 2, this again demonstrates how the key physics of ideally-stacked TMD heterostructures are very closely approximated by our simpler but asymmetric structures.
@@
Figure 4: $k_{z}$-mediated topological phases in asymmetric heterostructures (a) Band dispersions along an axis close to K-$\Gamma$-K taken with 81 eV photons for (from left to right) bulk WSe2, 3ML-WSe2, a (WSe2)3/(MoSe2)5 heterostructure, 3ML MoSe2 and a (MoSe2)3/(WSe2)3 heterostructure. Dashed rectangles show the energy-momentum region covered by the plots (b-d.)(b) Zoom ins of the spectra in (a) with enhanced colour contrast. (c) Top-layer projected DFT calculations for the equivalent regions (d) Equivalent plots with resolved chiral ($\langle S_{x}\rangle$) spin component. Fully-formed (i.e. ungapped) topological surface states (TSSs) are indicated in (b) and (c).
Though orbital-dependent, the presence of three-dimensional electronic behavior in these types of artificial heterostructures prompts the question of whether surface-localised electronic states can exist where global crystalline symmetries are truncated. The disparate electronic structures of top and bottom surfaces would have profound implications for any surface states, should they exist. Fortunately, TMD heterostructures are particularly well-suited to explore this eventuality as the presence of surface states in bulk TMDs is enforced by delocalised electron behavior along $k_{z}$ Bahramy et al. (2018); Clark et al. (2018, 2019).
@@
The families of three-fold symmetric 1T and 2H-structured TMDs possess a series of band inversions along their $k_{z}$ axes ($\parallel$ $\Gamma$-A) Huang et al. (2016); Yan et al. (2017); Noh et al. (2017); Li et al. (2017); Bahramy et al. (2018); Clark et al. (2018, 2019); Ghosh et al. (2019); Jiang et al. (2020); Mukherjee et al. (2020); Nurmamat et al. (2021); Clark et al. (2022); Chakraborty et al. (2023); Kimar et al. . These inversions, entirely contained to the chalcogen $p$-orbital manifold, generate up to five Dirac crossings (both bulk Dirac points and topological surface states (TSSs)) energetically stacked at the surface Brillouin centre. While the underpinning mechanism is equivalent for the dozen TMDs explicitly shown to host these states to date, the ladder of topological band crossings produced is delicately dependent on material-specific parameters, and thus are fingerprints of each individual system.
The minimal criteria to generate topological ladders in TMDs are only there exists a discrepancy in the $k_{z}$-bandwidths of in- and out-of-plane chalcogen $p$-orbital-derived states and $C_{3}$ symmetry is intact to maintain bulk Dirac crossings (though not crucial for the properties of surface Dirac states). The presence of bands well approximating a (discretised) $k_{z}$ dispersion along with orbital hopping asymmetries (discussed thus far for $d$-orbitals) imply that the above conditions are met in the experimental heterostructures considered here.
@@
Figure 4 overviews the realisation of non-trivial band topology in our proof-of-principle MoSe2-WSe2 heterostructures. In 2H-TMDs like WSe2, there is one experimentally resolvable TSS centred $\approx$ 1.9 eV below the VBM Bahramy et al. (2018), as observed in spectra from the bulk region of flake used for device construction (Fig.  4(a)). This is compared to band dispersions from the pair of WSe2-MoSe2 heterostructures and their respective 3ML top layers. Experimental zoom-ins of the region wherein the bulk band inversion occurs and top-layer projected DFT calculations are provided in Fig. 4(b) and (c), respectively. While quasi-linear band dispersions are present in all cases, both 3ML TMD flakes undergo significant band renormalization when stacked atop a dissimilar flake. In 3ML-WSe2, both experiment and theory demonstrate a small local SOC-mediated gap between a pair of bands at the binding energy where the Dirac crossing occurs in the bulk limit. When interfaced to form a 3ML-WSe2/5ML-MoSe2 heterostructure, however, the lower of these two bands shifts to higher binding energy to give way to a gapless TSS with well-defined upper and lower branches which merge into the surrounding bands at higher $k_{\parallel}$, as for the case of bulk WSe2. Similarly contrasting behavior is evident for the second, thinner, heterostructure: While the smaller atomic SOC strength of Mo precludes resolving the finite band gaps for the isolated 3ML flake, a distinct crossing point emerges when stacked, differing from the states of 3L or 5ML-MoSe2 (Sup. Fig. 1). With reference to Supplemental Figs. 6 and 7, one can verify that the spectral weight of the Dirac points are in each case localised to the top and bottom monolayers only, demonstrating the surface states are a property of the collective lattice unit and not its individual components, in line with the $k_{z}$-mediated topological phase transition from which they derive Bahramy et al. (2018).
@@
In contrast to $k_{z}$-mediated TSS found in naturally occurring TMDs, the appearances, energetic positions and characteristic chiral momentum-locked spin textures of these Dirac crossings (Fig. 4(d)), can differ between the two oppositely-orientated surfaces, with a $\approx 55$ meV discrepancy extracted for the specific (MoSe2)3/(WSe2)5 configuration from the layer-resolved calculations in Sup. Fig. 7, consistent with the experimental datasets from oppositely-orientated WSe2-MoSe2 structures in Fig. 4. This physics should apply to any pair of TMDs which are susceptible to $k_{z}$-mediated topological phase transitions, each with a unique set of Dirac crossings with surface-dependent behavior. For example, this could be utilized to tailor the five available Dirac cones in the 1T-structured TMD (semi)metals (Pd,Pt,Ni)(Se,Te)2, which occur near, or at, the Fermi level and have high sensitivity to relative hopping parameters in the unit cell Bahramy et al. (2018); Clark et al. (2019); Ferreira et al. (2021); Nicholson et al. (2021).

### v2

@@ III.2 Delocalised band formation in two-component stacks
To demonstrate how this concept of achieving bulk-like band characteristics with mixed material properties can be realised in real systems, Figure 2 overviews the electronic structure of a (WSe2)3/(MoSe2)5 (Fig. 2(a)) and a (MoSe2)3/(WSe2)3 (Fig. 2(e)) heterostructure constructed by interfacing two few-layer TMD flakes on graphite/n-type silicon substrates (see Methods). This pair of heterostructures thus closely approximate opposite surfaces of a single MoSe2-WSe2 system, and the modest total layer numbers enable unambiguous comparisons of stacked regions relative to the two component flakes. Photoemission spectra acquired from the stacked heterostructure regions are displayed in Fig. 2(b, f).
@@
As exemplified in Fig. 1(g-h), there is a one-to-one correspondence between the number of discrete sub-bands within the $d_{z^{2}}$-manifold at the $\Gamma$ point of 2H-structured TMDs and the number of monolayers within the system Chang et al. (2014); Le Févre et al. (2024). These bands collectively describe a single three-dimensional $d_{z^{2}}$-derived band which is quantised in $k_{z}$ due to the finite size of the system along the $c$-axis Chiang (2000); Milun et al. (2002); Le Févre et al. (2024). Therefore, as basic band theory stipulates that periodicity in momentum space cannot be achieved without periodicity in real space, the number of discrete $k_{z}$ sub-bands visible within the overlapped regions would not be expected to reflect the full extent of the heterostructures which lack a well-defined unit cell. Yet, the photoemission spectra in Fig. 2(b,f) demonstrate that the number of $k_{z}$ sub-bands is not only increased relative to equivalent datasets acquired from the constituent flakes (shown in full in Supplemental Fig. 1), but also matches the total layer number in real space (i.e. eight for (b) and six for (f)) of the entire heterostructure. The electron mean free path of less than 0.7 nm, applicable to our photoemission experiments, precludes that the experimental result originates from a superposition of spectra from weakly-interacting top and bottom TMD flakes which each have thicknesses of $\sim$ 0.7 nm per monolayer Pan and Zahn (2022). This alternate scenario is further discarded through constant-energy $k_{x}$-$k_{y}$ contour mapping in Supplemental Figure 2. Instead, the six or eight distinct $d_{z^{2}}$ sub-bands at $\Gamma$ are delocalised throughout the structure, and are therefore accessible to a surface sensitive probe.
@@
While counterintuitive, these findings are entirely consistent with the discussion surrounding Figure 1: The array of hybridized bands at $\Gamma$ belonging to the collective band structure of the finite six or eight layer crystal structure closely mimics the discretized $k_{z}$-projected band manifold of ordered systems. For bands derived from orbitals with significant interlayer hopping potential, the $k_{z}$ dispersion is therefore effectively constructed without the need for a periodic repeating unit. To further test this description, the experimental band dispersions of both WSe2-MoSe2 heterostructures are compared to layer-projected, density functional theory (DFT)-based slab calculations in Fig. 2(c,g), (see Methods). The electronic structure around the $\Gamma$ point is extremely well reproduced in each case, verifying the present interpretation while also demonstrating a negligible role of collective electron phases and large twist angles (the latter chosen experimentally in order to decouple Moiré effects, see Sup. Fig. 3), which are both neglected by these calculations.
@@
To further scrutinize the observed three-dimensional behavior, in Figure 2(d,h) we track the evolution of the delocalised states at $\Gamma$ with changing photon energy, thus mapping their behavior with changing $k_{z}$ Damascelli (2004). These spectra can be compared to equivalent datasets for a three-layer MoSe2 and WSe2 flake (Sup. Fig. 1 (f,i)). For both the individual TMD flakes and the artificial heterostructures, each $d_{z^{2}}$-subband is two-dimensional (i.e. non-dispersive along $k_{z}$) in-line with their assignment as dimensionally confined states, though each exhibits a photon energy-dependent spectral weight. In the the eight- or six-layer hybrid systems (Fig. 2(d,h)), it is apparent that these spectral weight variations are periodic in $k_{z}$, with each sub-band illuminated sequentially. This is consistent with their collective assignment as a singular $d_{z^{2}}$-orbital derived band, which remains quantized in $k_{z}$ due to the finite length scales of the total hybrid systems along the asymmetric $c$-axis. Though discrete, this periodic behavior is exactly as would be expected for three-dimensional bands in a bulk 2H-structured TMD system. Indeed, the overlaid solid lines, which well capture the observed periodicity, are a tight-binding model for an idealized bulk structure with alternating WSe2 and MoSe2 layers along the $c$-axis. We note that, due to significant $k_{z}$ broadening in photoemission generally Damascelli (2004), the majority of these sub bands can be captured simultaneously when probed with a single photon energy, as evidenced by the in-plane band dispersions shown in Fig. 2(b,f).
In Supplemental Figure 4, we show how equivalent phenomena is observed for mismatched combinations of WSe2 and MoSe2 bilayers: the number of $k_{z}$ sub-bands within the $d_{z^{2}}$-manifold reflects the full $c$-axis extent of the system despite ill-defined unit cells, again demonstrating how the periodic $k_{z}$-dispersions expected for ordered solids are emulated when interfacing two-or-more TMD flakes in the intermediate layer regime. We stress that, while the observed three-dimensional band dispersions throughout this Section exhibit finite-size effects, the key properties of bulk TMDs are, in general, present and fully-developed in their six (or fewer)-monolayer variants despite $k_{z}$ quantisation Kim and Choi (2021); Hlevyack et al. (2021); Deng et al. (2019); Lan et al. (2021); Fridman et al. (2025); Chang et al. (2014); Le Févre et al. (2024). Nevertheless, we expect that the underpinning mechanism driving delocalised states along $k_{z}$ to hold for pairings of TMD flakes with larger total thicknesses, thus producing artificial 3D band properties closer to the true bulk limit.
III.3 Orbital-dependent band localisation
@@
While the presence of delocalised electron behavior within the $d_{z^{2}}$ orbital manifolds of the finite heterostructures is clear, the same is not true for regions of the electronic structure derived from in-plane orbitals. This stark contrast in orbital-dependent coupling is a hallmark of 2H-type transition metal dichalcogenides generally. The details of the $d_{\parallel}$ K-point valence bands are largely equivalent in monolayer and bulk 2H-TMD systems due to their minimal out-of-plane hopping potential Riley et al. (2015); Mak et al. (2010); Xiao et al. (2012), whereas $d_{z^{2}}$-derived $\Gamma$-point states exhibit substantial out-of-plane coupling to drive significant $k_{z}$ dispersion in bulk materials  Riley et al. (2014); Latzke et al. (2015); Kim et al. (2016); Alarab et al. (2023); Le Févre et al. (2024); Chang et al. (2014), thus underpinning the indirect-to-direct bandgap transition observed in few-layer systems Mak et al. (2010); Sun et al. (2016); Ernandes et al. (2021). This natural orbital disparity interplays with the inherent asymmetry of heterostructures constructed by interfacing two dissimilar TMD flakes, ensuring a spatial dependence to the broader electronic structure along the $c$-axis between limits defined by the distinct component materials. The scale of this dependence is therefore dependent on orbital type.
@@
In Figure 3 and Supplemental Figs. 5-6, this real space evolution is explored in full. In Fig. 3(a), the orbitally-resolved WSe2 and MoSe2 terminations of the experimentally realized (MoSe2)3/(WSe2)3 heterostructure are compared. The valence band maxima at K and $\Gamma$, derived exclusively from $d_{\parallel}$ and $d_{z^{2}}$-orbitals respectively, represent the two extremes of $c$-axis delocalisation: The details of the spin-orbit splitting at K is determined only by the local crystal structure, while the band structure at $\Gamma$ is equivalent between terminations and indeed between the top and bottom layers (shown explicitly in Supplemental Figure 5). The rest of the band structure behaves similarly, with the real-space electronic structure dependence directly determined by the $d_{\parallel}$/ $d_{z^{2}}$ orbital ratio, further influenced by hybridization between bands of differing orbital origins. In artificial, asymmetric heterostructures, this therefore produces an overall layer-dependent electronic structure despite the presence of fully delocalised band features. This is in stark contrast to naturally occurring TMDs, wherein the overall symmetry of the system precludes real space band structure variations despite a similar orbital-dependent localisation.
@@
The asymmetry of our artificial heterostructures can also be expected to modify the spin structure relative to natural TMD systems. As discussed earlier, in a natural 2H-TMD, each constituent TMD layer, owing to its $D_{3h}$ symmetry, can only exhibit an out-of-plane spin-orbit field, whose direction is opposite to that of the neighbouring layers. Consequently, this implies that in-plane spin textures, including chiral spin-momentum locking types arising from the Rashba effect Bychkov and Rashba (1984), are effectively suppressed. Furthermore, the hidden out-of-plane spin textures of each layer can only become accessible at the surface, where translational symmetry breaking disrupts the natural cancellation of $B_{\text{SO}}$ along the stacking axis.
@@
In contrast, our artificial structures can naturally accommodate both in-plane Rashba-type and out-of-plane Zeeman-type spin textures due to their unique heterostructural properties. Specifically, the chiral Rashba spin polarisation can emerge naturally from the potential gradient formed at the MoSe2/WSe2 interface and propagate throughout the entire structure via $d_{z^{2}}$ and other out-of-plane orbitals. To elucidate the implications of this, in Fig. 3(b-f), we compare a (MoSe2)3/(WSe2)3 heterostructure to an idealised six-layer (MoSe2/WSe2)3 system. The calculated layer-projected spin polarisation for the chiral ($x$) and out-of-plane ($z$) components is shown for both the asymmetric and idealised cases. As can be observed, in both configurations, strikingly similar patterns of in-plane and out-of-plane spin polarisation emerge, which remain valley-specific. Particularly, in both cases, the chiral ($x$) component is significantly enhanced around the $\Gamma$ point, where the $d_{z^{2}}$ contribution is dominant. As with the discussions surrounding the $k_{z}$ dispersions on show in Fig. 2, this again demonstrates how the key physics of ideally-stacked TMD heterostructures are very closely approximated by our simpler but asymmetric structures.
III.4 Creation of layer-dependent semimetallic phases
Figure 4: Local semimetallic fermiology in an asymmetric WSe2-NbSe2 heterostructure (a) Optical image of the heterostructure region. The few-monolayer thick top and bottom layers are highlighted in red and green respectively. The overlap region of interest (ROI) is highlighted in purple. (b) Band dispersions of the ROI along axes close to the $\Gamma-K$ direction. The inset is a $k_{x}$-$k_{y}$ contour at the Fermi level, taken with 90 eV photons. (c) Density functional theory based slab calculations, projected through the full heterostructure (left) and to the top layer (right) along the $\Gamma-K$ direction. (d) $k_{z}$ dispersion (hvv=60-144 eV) at $k_{\parallel}=0$. Dashed lines indicate the slices in (b), and solid lines are a tight-binding model (TBM) of bulk WSe2 offset in energy to match the experimental scenario. An inner potential of 16 eV is chosen to optimize the match to the periodicity of the calculated band structure. $k_{z}$ sub-bands of the $d_{z^{2}}$ valence band maxima are numbered from low binding energy to high. We note that the spectrum in (b) is partially superimposed with that from pristine 4L-WSe2, most likely due to the less-than-pristine overlap region. We note that, due to air sensitivity of air-exposed NbSe2, atomic force microscopy (AFM) was employed to determine the thickness of the NbSe2 bottom layer instead of in situ photoemission. The step height from the WSe2 regions (red outline in (a)) to the overlap region (purple) was found to be $\sim$2.25 nm consistent with 2 ML NbSe2 Wang et al. (2017)
The asymmetric WSe2-MoSe2 heterostructures discussed thus far exhibit dissimilar top and bottom surfaces due to orbital-dependent band localisation. While the global symmetries of the asymmetric unit are fundamentally altered relative to naturally occurring TMDs or to an idealized stack, its layer-dependent electronic structure is derivative of the bulk forms of the component materials. For the specific case considered thus far, the similarities in the WSe2 and MoSe2 lattice constants Kang et al. (2013) and in the energetics of the relevant W 5dd and Mo 4dd orbital manifolds precludes significant deviations of the formed hybrid electronic structure from those of the parent compounds. We stress however, that such a similarity in material components is not a necessary criterion to stabilize dimensionally-mixed frameworks: The mechanism underpinning our findings should apply generally to pairs of 2D materials with non-negligible out-of-plane hopping, even when the two components differ substantially. Such a mismatching could be used to engineer entirely unique quantum phases.
To exemplify the power of this eventuality, in Fig. 4, we showcase a (WSe2)4/(NbSe2)2 heterostructure. NbSe2 is metallic, whereas WSe2 is semiconducting. Consequently, the energy bands in WSe2 are effectively lower than those in NbSe2 by the energy gained from filling the valence bands with an additional electron, transferred from the W+4 ion compared to the Nb+3 ion. Despite this, not only do the number of $k_{z}$ subbands at $\Gamma$ continue to reflect the total layer number of the heterostructure thereby demonstrating the presence of fully delocalised bands, but the system becomes semimetallic, with the shallowest binding energy sub-band at the $\Gamma$ point forming an Fermi surface, shown inset in Fig. 4b and in Supplemental Fig. 7. This Fermi pocket at $\Gamma$ is the only Fermi band crossing on the top layers of the artificial heterostructure, unlike the electronic structure of any naturally occurring bulk or few-layer TMD. In Fig. 4(c), the experimental picture is compared to DFT calculations. While, as with the previous heterostructures, there is strong quantitative agreement, we note the presence of additional fine structure to the bands near the Fermi level at $\Gamma$ relative to the equivalent bands in the previous heterostructures. We attribute this to hybridisation effects between the W $d_{z}^{2}$ manifold and the Se $p_{z}$ manifold from NbSe2, which overlap in this hybrid heterostructure. Nevertheless, as verified by the layer-resolved calculations in Sup. Fig. 8, all sub-bands at $\Gamma$ are again present and at the same energy in each layer of the heterostructure, evidencing the presence of fully delocalised electrons along the $c$-axis and hence the formation of a discretized three-dimensional band manifold. Indeed, the $k_{z}$ periodicity in the spectral weight of the experimentally observed sub-bands, shown in Fig. 4(d), is again well matched to the periodicity of bulk TMD calculations.
The observed $c$-axis delocalisation, persisting even for this energetically-mismatched scenario, is facilitated by the lowering of the total energy of the system by partially lifting the constraints of 2D quantization. Along the $c$-axis, electrons form standing waves due to the finite thickness of the artificial crystals. This effect can be expected to vanish as one or both flakes approach the infinite-layer limit and electrons at the interface gain access to a bulk reservoir, ultimately preserving the quantized electronic structure of the remaining few-layer component. This is tested and shown explicitly in Supplemental Figure 9 through the characterization of a second overlap region within the sample in Fig. 4(a); between a 4ML-WSe2 flake and a many-layer NbSe2 crystal. While charge transfer from the semiconducting top-layer still occurs in addition to some normalization of band features relative to the isolated 4ML-WSe2 top-layer, there are no additional well-defined $d_{z^{2}}$-subbands.
III.5 $k_{z}$-mediated topological phases
@@
Figure 5: $k_{z}$-mediated topological phases in asymmetric heterostructures (a) Band dispersions along an axis close to $\Gamma-K$ taken with 81 eV photons for (from left to right) bulk WSe2, 3ML-WSe2, a (WSe2)3/(MoSe2)5 heterostructure, 3ML MoSe2 and a (MoSe2)3/(WSe2)3 heterostructure. Dashed rectangles show the energy-momentum region covered by the plots (b-d.)(b) Zoom ins of the spectra in (a) with enhanced colour contrast. (c) Top-layer projected DFT calculations for the equivalent regions (d) Equivalent plots with resolved chiral ($\langle S_{x}\rangle$) spin component. Fully-formed (i.e. ungapped) topological surface states (TSSs) are indicated in (b) and (c).
Though orbital-dependent, the presence of delocalised electronic behavior in asymmetric artificial heterostructures prompts the question of whether additional surface-localised states are produced by the truncation of global crystalline symmetries. While surface states, both topologically trivial and non-trivial, are commonplace in naturally occurring bulk TMDs, the disparate surfaces unique to asymmetric heterostructures would result in pairs of surface states with mismatched, surface-specific, properties.
@@
To explore this potential, in Fig. 5 we search for surface states enforced by topological phase transitions driven by out-of-plane delocalised electron behavior. Bahramy et al. (2018); Clark et al. (2018, 2019). The families of three-fold symmetric 1T and 2H-structured TMDs possess a series of band inversions along their $k_{z}$ axes ($\parallel$ $\Gamma$-A) Huang et al. (2016); Yan et al. (2017); Noh et al. (2017); Li et al. (2017); Bahramy et al. (2018); Clark et al. (2018, 2019); Ghosh et al. (2019); Jiang et al. (2020); Mukherjee et al. (2020); Nurmamat et al. (2021); Clark et al. (2022); Chakraborty et al. (2023); Kimar et al. . These inversions, entirely contained to the chalcogen $p$-orbital manifold, generate up to five Dirac crossings (both bulk Dirac points and topological surface states (TSSs)) energetically stacked at the surface Brillouin centre. While the underpinning mechanism is equivalent for the dozen TMDs explicitly shown to host these states to date, the ladder of topological band crossings produced is delicately dependent on material-specific parameters, and thus are fingerprints of each individual system. The minimal criteria to generate topological ladders in TMDs are only that there exists a discrepancy in the $k_{z}$-bandwidths of in- and out-of-plane chalcogen $p$-orbital-derived states, and that $C_{3}$ symmetry is intact to maintain bulk Dirac crossings (though not crucial for the properties of surface Dirac states). The presence of bands well approximating a (discretized) $k_{z}$ dispersion along with orbital hopping asymmetries (discussed thus far for $d$-orbitals) imply that the above conditions are met in the experimental heterostructures considered here. Previous ARPES studies have shown that 4-5 monolayers is sufficient to stabilize the topologically non-trivial properties of bulk TMD systems in full Hlevyack et al. (2021); Deng et al. (2019). Such physics is therefore relevant to the intermediate-layer regime considered here.
@@
Figure 5 overviews the realisation of non-trivial band topology in our proof-of-principle MoSe2-WSe2 heterostructures. In bulk 2H-TMDs like WSe2, there is one experimentally resolvable TSS centred $\sim$ 1.9 eV below the VBM Bahramy et al. (2018), as observed in spectra from the bulk region of the flake used for device construction (Fig.  5(a)). This is compared to band dispersions from the pair of WSe2-MoSe2 heterostructures and their respective 3ML top layers. Experimental zoom-ins of the region wherein the bulk band inversion occurs and top-layer projected DFT calculations are provided in Fig. 5(b) and (c), respectively. While quasi-linear band dispersions are present in all cases, both 3ML TMD flakes undergo significant band renormalization when stacked atop a dissimilar flake. In 3ML-WSe2, both experiment and theory demonstrate a small local SOC-mediated gap between a pair of bands at the binding energy where the Dirac crossing occurs in the bulk limit. When interfaced to form a 3ML-WSe2/5ML-MoSe2 heterostructure, however, the lower of these two bands shifts to higher binding energy to give way to a gapless TSS with well-defined upper and lower branches which merge into the surrounding bands at higher $k_{\parallel}$, as for the case of bulk WSe2. Similarly contrasting behavior is evident for the second, thinner, heterostructure: While the smaller atomic SOC strength of Mo precludes resolving the finite band gaps for the isolated 3ML flake, a distinct crossing point emerges when stacked, differing from the states of 3L or 5ML-MoSe2 (Sup. Fig. 1). With reference to Supplemental Figs. 5 and 6, one can verify that the spectral weight of the Dirac points are in each case localised to the top and bottom monolayers only, demonstrating the surface states are a property of the collective lattice unit and not its individual components, in line with the $k_{z}$-mediated topological phase transition from which they derive Bahramy et al. (2018).
@@
In contrast to $k_{z}$-mediated TSS found in naturally occurring TMDs, the appearances, energetic positions and characteristic chiral momentum-locked spin textures of these Dirac crossings (Fig. 5(d)), can differ between the two oppositely-orientated surfaces, with a $\sim 55$ meV discrepancy extracted for the specific (MoSe2)3/(WSe2)5 configuration from the layer-resolved calculations in Sup. Fig. 6, consistent with the experimental datasets from oppositely-orientated WSe2-MoSe2 structures in Fig. 5. This physics should apply to any pair of TMDs which are susceptible to $k_{z}$-mediated topological phase transitions, each with a unique set of Dirac crossings with surface-dependent behavior. This could be utilized to drastically augment the five available Dirac cones in the 1T-structured TMD (semi)metals (Pd,Pt,Ni)(Se,Te)2 which have high sensitivity to relative hopping parameters in the unit cell Bahramy et al. (2018); Clark et al. (2019); Ferreira et al. (2021); Nicholson et al. (2021). For example, compared to PtSe2, PdTe2 hosts an additional $k_{z}$-band crossing to produce a TSS at Fermi level Clark et al. (2018). It follows that an asymmetric heterostructure formed from these two compounds would host a hybrid $k_{z}$-dispersion with a band inversion close to the quantum critical point, potentially enabling in situ switching of TSS contributions to transport.
IV CONCLUSIONS